# Spectrum-Aware Debiasing: A Modern Inference Framework with Applications to Principal Components Regression

### v4

@@ B.3.1 Oracle VAMP proofs
This then implies that $\delta_{*}$ is the unique fixed point of $g(\cdot)$ over $[0,\delta_{*}]$, and $\lim_{t\to\infty}g^{(t)}(\delta)=\delta_{*}$ for any $\delta\in[0,\delta_{*}]$. Observe from (80) that $\delta_{1t}=\delta_{12}$ for all $t\geq 2$, so $\delta_{t,t+s}=g^{(t-1)}(\delta_{1,1+s})=g^{(t-1)}(\delta_{12})$ for any $s\geq 1$. Then $\lim_{\min(s,t)\to\infty}\delta_{st}=\delta_{*}$ follows.
It remains to show (81). Using Proposition B.1,
@@
using Proposition B.6. The convergence of iterates $\mathbf{r}_{1t},\mathbf{r}_{2t}$ follows from $\mathbf{r}_{2t}=\mathbf{x}^{t}+\bm{\beta}^{\star},\mathbf{r}_{1t}=\mathbf{y}^{t}+\bm{\beta}^{\star}+\mathbf{e}$. The convergence of $\hat{\mathbf{x}}_{1t},\hat{\mathbf{x}}_{2t}$ follows from the fact they can be expressed as Lipschitz function applied to iterates $\mathbf{r}_{1,t-1}$ and $\mathbf{r}_{2t}$, i.e. (59a) and (59c). ∎
B.3.2 Track regularized estimator using VAMP iterates
Proof of Proposition B.8.
We assume that $c_{0}>0$. Proof for the case where $d_{-}>0$ is completely analogous. From strong convexity of the penalty function, almost surely, for all sufficiently large $p$,
holds for any $\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right)\in\mathbf{X}^{\top}\left(\mathbf{X}\hat{\mathbf{x}}_{1t}-\mathbf{y}\right)+\partial h\left(\hat{\mathbf{x}}_{1t}\right)$ where $\partial h$ denotes sub-gradients of $h$. So we can take $\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right)=\mathbf{X}^{\top}\left(\mathbf{X}\hat{\mathbf{x}}_{1t}-\mathbf{y}\right)+\gamma_{*}\left(\mathbf{r}_{1,t-1}-\hat{\mathbf{x}}_{1t}\right)$ because $\hat{\mathbf{x}}_{1t}=\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathbf{r}_{1,t-1}\right)\Leftrightarrow\mathbf{r}_{1,t-1}-\hat{\mathbf{x}}_{1t}\in\gamma_{*}^{-1}\partial h\left(\hat{\mathbf{x}}_{1t}\right).$ By Cauchy-Schwartz inequality, we have that
@@ B.3.2 Track regularized estimator using VAMP iterates
=$\stackrel{{\scriptstyle(a)}}{{=}}\left(1-\frac{\gamma_{*}}{\eta_{*}}\right)\left(\mathbf{X}^{\top}\mathbf{X}+\gamma_{*}\mathbf{I}_{p}\right)\left(\mathbf{r}_{2t}-\mathbf{r}_{2,t-1}\right)+\left(\mathbf{X}^{\top}\mathbf{X}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{I}_{p}\right)\hat{\mathbf{x}}_{2,t-1}$
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}\left(1-\frac{\gamma_{*}}{\eta_{*}}\right)\left(\mathbf{X}^{\top}\mathbf{X}+\gamma_{*}\mathbf{I}_{p}\right)\left(\mathbf{r}_{2t}-\mathbf{r}_{2,t-1}\right)$
@@
which follows from (59c). We thus have that almost surely
@@
which along with (88) implies that
@@
By (89), we also have that
@@
B.4 Supporting proofs for result B
B.4.1 Properties of sample adjustment equation
Proof of Lemma B.9.
@@ B.4.1 Properties of sample adjustment equation
[…]
@@
$01-\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}\frac{\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}-\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}1>$ 0 (93)
@@
B.4.2 Population limit of the adjustment equation
Proof of Lemma B.11.
@@ B.4.2 Population limit of the adjustment equation
$01-\mathbb{E}\frac{\gamma\mathbb{I}(\mathsf{U}\neq+\infty,0)}{\gamma+\mathsf{U}}-\mathbb{P}(\mathsf{U}=0)>$ 0 (96)
@@
Note that $\mathbb{P}(\mathsf{U}\neq 0)>0$ if and only if $\mathbb{P}(\mathsf{U}\neq 0,+\infty)>0$ or $\mathbb{P}(\mathsf{U}=+\infty)>0$. This shows that (96) holds and thus $g_{\infty}$ is well-defined since $\mathbb{P}(\mathsf{U}\neq 0)>0$ by Lemma A.10.
It follows from (20a), (20c) and (16) that $\gamma_{*}$ is a solution of the equation $g_{\infty}(\gamma)=1$ We prove that $\gamma_{*}$ is a unique solution by showing $g_{\infty}$ is strictly increasing. Applying [90, Proposition A.2.1], we obtain that $g_{\infty}$ is differentiable and can be differentiated inside the expectation as follows
@@
To prove $g_{\infty}^{\prime}(\gamma)>0,\forall\gamma\in(0,+\infty)$, note that if $\mathbb{P}(\mathsf{U}\neq+\infty,0)>0$, then $\mathbb{E}\frac{\mathsf{U}\mathbb{I}(\mathsf{U}\neq+\infty,0)}{(\gamma+\mathsf{U})^{2}}>0$ and the above will be positive. Also note that if $\mathbb{P}(\mathsf{U}=0)>0$, then $\mathbb{I}\left(\mathsf{D}^{2}>0\right)\left(\mathsf{D}^{2}\frac{1}{\gamma^{2}}\right)\mathbb{P}(\mathsf{U}=0)>$ 0 with positive probability and the above will be positive. Note that $\mathbb{P}(\mathsf{U}\neq+\infty)>0$ if and only if $\mathbb{P}(\mathsf{U}\neq 0$ and $\mathsf{U}\neq+\infty)>0$ or $\mathbb{P}(\mathsf{U}=0)>0$. Therefore, the positivity of $g_{\infty}^{\prime}(\gamma)$ follows from $\mathbb{P}(\mathsf{U}\neq+\infty)>0$ which holds by Lemma A.10. The proof is now complete. ∎
Proof of Proposition B.13.
@@
This follows from $\mathbf{r}_{*}\in\hat{\bm{\beta}}+\frac{1}{\gamma_{*}}\partial h(\hat{\bm{\beta}})$ and the equivalence relation $\mathbf{r}_{*}\in\hat{\bm{\beta}}+\frac{1}{\gamma_{*}}\partial h(\hat{\bm{\beta}})\Leftrightarrow\hat{\bm{\beta}}=\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathbf{r}_{*}\right)$. The former is a consequence of the KKT condition $\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})\in\partial h(\hat{\bm{\beta}})$ and the latter follows from Proposition A.7, $(a)$. Also note that for any $\gamma>0$,
@@
$\lim_{p\rightarrow\infty}\frac{1}{p}\sum_{i=1}^{p}\frac{1}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{i}\right)}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\lim_{p\rightarrow\infty}\frac{1}{p}\sum_{j=1}^{p}\frac{1}{\gamma+h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(r_{*,j}\right)\right)}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{E}\frac{1}{\gamma+h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\right)}$
where $(a)$ follows from (97) and $(b)$ follows from Theorem 5.1, Proposition A.5 and (98). This, along with Assumption 2 and Proposition A.4, implies that almost surely
@@
Almost sure convergence (72) follows from (99), Proposition A.5 and the fact from Lemma B.11 that $01+\left(\mathsf{D}^{2}-\gamma\right)\mathbb{E}\frac{1}{\gamma+\mathsf{U}}>$ 0 almost surely. ∎
B.5 Finite or single coordinate inference under exchangeability
@@
We only show that $\frac{{r}_{*,i}-{\beta}^{\star}_{i}}{\sqrt{\tau_{*}}}\Rightarrow N(0,1)$ for $\mathbf{r}_{*}$ defined in (34). (27) then follows from consistency of $\hat{\tau}_{*}$ and $\widehat{\mathsf{adj}}$ (cf. Theorem 5.2) and the Slutsky’s theorem. Let $\mathbf{U}\in\mathbb{R}^{p\times p}$ denote a permutation operator drawn uniformly at random independent of $\bm{\beta}^{\star},\mathbf{X},\bm{\varepsilon}$. We have that
@@
where we use =$\stackrel{{\scriptstyle L}}{{=}}$ to denote equality in law. Note that
@@
[…]
@@
Below we prove the Corollary for =$\mathcal{L}=\{i,k\},i\neq k$. The general case is analogous. For standard basis $\mathbf{e}_{i},\mathbf{e}_{k}$, and any constant $c_{1},c_{2}\in\mathbb{R}$,
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{P}\left(\frac{\mathbf{e}_{i}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{i}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{1},\frac{\mathbf{e}_{k}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{k}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{2}\right)$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{E}\left(\mathbb{P}\left(\frac{\mathbf{e}_{i}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{i}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{1},\frac{\mathbf{e}_{k}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{k}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{2}\mid\mathcal{F}\left(\bm{\beta}^{\star},\bm{\varepsilon},\mathbf{X}\right)\right)\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\mathbb{E}\frac{1}{p(p-1)}\sum_{j_{1}\neq j_{2}\in[p]}\mathbb{I}\left(\frac{1}{\sqrt{\tau_{*}}}\left(r_{*,j_{1}}-\beta_{j_{1}}^{\star}\right)<c_{1}\right)\mathbb{I}\left(\frac{1}{\sqrt{\tau_{*}}}\left(r_{*,j_{2}}-\beta_{j_{2}}^{\star}\right)<c_{2}\right)$
@@
where $\mathsf{Z}\sim N(0,1)$. Here, we used Theorem 5.2 and Proposition A.5. Using dominated convergence theorem, we conclude that
@@
B.6 Hypothesis testing and confidence intervals

### v6

@@ C.5.1 Oracle VAMP proofs
This then implies that $\delta_{*}$ is the unique fixed point of $g(\cdot)$ over $0,\delta_{*}]$, and $\lim_{t\to\infty}g^{(t)}(\delta)=\delta_{*}$ for any $\delta\in[0,\delta_{*}]$. Observe from (94) that $\delta_{1t}=\delta_{12}$ for all $t\geq 2$, so $\delta_{t,t+s}=g^{(t-1)}(\delta_{1,1+s})=g^{(t-1)}(\delta_{12})$ for any $s\geq 1$. Then $\lim_{\min(s,t)\to\infty}\delta_{st}=\delta_{*}$ follows.
It remains to show (95). Using Proposition C.3,
@@
using Proposition C.8. The convergence of iterates $\mathbf{r}_{1t},\mathbf{r}_{2t}$ follows from $\mathbf{r}_{2t}=\mathbf{x}^{t}+\bm{\beta}^{\star},\mathbf{r}_{1t}=\mathbf{y}^{t}+\bm{\beta}^{\star}+\mathbf{e}$. The convergence of $\hat{\mathbf{x}}_{1t},\hat{\mathbf{x}}_{2t}$ follows from the fact they can be expressed as Lipschitz function applied to iterates $\mathbf{r}_{1,t-1}$ and $\mathbf{r}_{2t}$, i.e. (72a) and (72c). ∎
C.5.2 Track regularized estimator using VAMP iterates
Let us first prove the following lemma
Lemma C.16.
Recall the objective function $\mathcal{L}$ defined in (1). The vector
$\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right):=\mathbf{X}^{\top}\left(\mathbf{X}\hat{\mathbf{x}}_{1t}-\mathbf{y}\right)+\gamma_{*}\left(\mathbf{r}_{1,t-1}-\hat{\mathbf{x}}_{1t}\right)$
is a subgraident of $\mathcal{L}$ at $\hat{\mathbf{x}}_{1t}$. We also have that almost surely,
$\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right)\right\|_{2}^{2}=0.$
Proof of Lemma C.16.
Let $\partial h$ denotes sub-gradients of $h$. We have that
$\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right)=\mathbf{X}^{\top}\left(\mathbf{X}\hat{\mathbf{x}}_{1t}-\mathbf{y}\right)+\gamma_{*}\left(\mathbf{r}_{1,t-1}-\hat{\mathbf{x}}_{1t}\right)\in\mathbf{X}^{\top}\left(\mathbf{X}\hat{\mathbf{x}}_{1t}-\mathbf{y}\right)+\partial h\left(\hat{\mathbf{x}}_{1t}\right)$
because
$\hat{\mathbf{x}}_{1t}=\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathbf{r}_{1,t-1}\right)\Leftrightarrow\mathbf{r}_{1,t-1}-\hat{\mathbf{x}}_{1t}\in\gamma_{*}^{-1}\partial h\left(\hat{\mathbf{x}}_{1t}\right).$
@@ C.5.2 Track regularized estimator using VAMP iterates
$\stackrel{{\scriptstyle(a)}}{{=}}\left(1-\frac{\gamma_{*}}{\eta_{*}}\right)\left(\mathbf{X}^{\top}\mathbf{X}+\gamma_{*}\mathbf{I}_{p}\right)\left(\mathbf{r}_{2t}-\mathbf{r}_{2,t-1}\right)+\left(\mathbf{X}^{\top}\mathbf{X}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{I}_{p}\right)\hat{\mathbf{x}}_{2,t-1}$
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\left(1-\frac{\gamma_{*}}{\eta_{*}}\right)\left(\mathbf{X}^{\top}\mathbf{X}+\gamma_{*}\mathbf{I}_{p}\right)\left(\mathbf{r}_{2t}-\mathbf{r}_{2,t-1}\right)$
@@
which follows from (72c). It then follows from Proposition C.8 that almost surely
@@
as required.
Proof of Proposition C.10.
Let us first consider the case $c_{0}>0$ from Assumption 4. From strong convexity of the penalty function, almost surely, for all sufficiently large $p$,
where $\mathcal{L}^{\prime}\left(\hat{\mathbf{x}}_{1t}\right)$ is the subgradient of $\mathcal{L}$ defined in Lemma C.16.
By Cauchy-Schwartz inequality, we have that
which along with Lemma C.16 implies that
@@
By (101) and Proposition C.8, we also have that
@@
The proof for the other case in Assumption 4—that is, when $\lim_{p\to\infty}\min_{i\in[p]}d_{i}^{2}>c_{1}$—is almost identical; the only difference is that for all sufficiently large $p$, (102) and (103) hold with $c_{0}$ replaced by $c_{1}$. ∎
C.6 Supporting proofs for result B
C.6.1 Properties of sample adjustment equation
Proof of Lemma C.11.
@@ C.6.1 Properties of sample adjustment equation
$\begin{aligned} g_{p}(\gamma)=&\frac{1}{p}\sum_{i:d_{i}\neq 0}\frac{1}{\frac{1}{p}\left(\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq+\infty,0}\frac{d_{i}^{2}-\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}+\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}\frac{d_{i}^{2}-\gamma}{\gamma}\right)+1}\\ &+\frac{1}{p}\sum_{i:d_{i}=0}\frac{1}{\frac{1}{p}\left(\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}\frac{-\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}-\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}1\right)+1}\end{aligned}.$ (106)
@@
$1-\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)\neq 0,+\infty}\frac{\gamma}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)}-\frac{1}{p}\sum_{j:h^{\prime\prime}\left(\hat{{\beta}}_{j}\right)=0}1>0$ (107)
@@
C.6.2 Population limit of the adjustment equation
Proof of Lemma C.13.
@@ C.6.2 Population limit of the adjustment equation
$1-\mathbb{E}\frac{\gamma\mathbb{I}(\mathsf{U}\neq+\infty,0)}{\gamma+\mathsf{U}}-\mathbb{P}(\mathsf{U}=0)>0$ (110)
@@
Note that $\mathbb{P}(\mathsf{U}\neq 0)>0$ if and only if $\mathbb{P}(\mathsf{U}\neq 0,+\infty)>0$ or $\mathbb{P}(\mathsf{U}=+\infty)>0$. This shows that (110) holds and thus $g_{\infty}$ is well-defined since $\mathbb{P}(\mathsf{U}\neq 0)>0$ by Lemma B.1.
It follows from (20a), (20c) and (16) that $\gamma_{*}$ is a solution of the equation $g_{\infty}(\gamma)=1$ . We prove that $\gamma_{*}$ is a unique solution by showing $g_{\infty}$ is strictly increasing. Applying [105, Proposition A.2.1], we obtain that $g_{\infty}$ is differentiable and can be differentiated inside the expectation as follows
@@
To prove $g_{\infty}^{\prime}(\gamma)>0,\forall\gamma\in(0,+\infty)$, note that if $\mathbb{P}(\mathsf{U}\neq+\infty,0)>0$, then $\mathbb{E}\frac{\mathsf{U}\mathbb{I}(\mathsf{U}\neq+\infty,0)}{(\gamma+\mathsf{U})^{2}}>0$ and the above will be positive. Also note that if $\mathbb{P}(\mathsf{U}=0)>0$, then $\mathbb{I}\left(\mathsf{D}^{2}>0\right)\left(\mathsf{D}^{2}\frac{1}{\gamma^{2}}\right)\mathbb{P}(\mathsf{U}=0)>$ 0 with positive probability and the above will be positive. Note that $\mathbb{P}(\mathsf{U}\neq+\infty)>0$ if and only if $\mathbb{P}(\mathsf{U}\neq 0$ and $\mathsf{U}\neq+\infty)>0$ or $\mathbb{P}(\mathsf{U}=0)>0$. Therefore, the positivity of $g_{\infty}^{\prime}(\gamma)$ follows from $\mathbb{P}(\mathsf{U}\neq+\infty)>0$ which holds by Lemma B.1. The proof is now complete. ∎
Proof of Proposition C.15.
@@
[…]
@@
$\lim_{p\rightarrow\infty}\frac{1}{p}\sum_{i=1}^{p}\frac{1}{\gamma+h^{\prime\prime}\left(\hat{{\beta}}_{i}\right)}$ $\stackrel{{\scriptstyle(a)}}{{=}}\lim_{p\rightarrow\infty}\frac{1}{p}\sum_{j=1}^{p}\frac{1}{\gamma+h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(r_{*,j}\right)\right)}$ (113)
$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{E}\frac{1}{\gamma+h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\right)}$
where $(a)$ follows from (111) and $(b)$ follows from Theorem C.1, Proposition A.5 and (112). An immediate consequence is that almost surely for all sufficiently large $p$, there must exist some $i\in[p]$ such that $h^{\prime\prime}(\hat{\beta}_{i})\neq+\infty$. This is because the RHS is bounded away from 0 for any fixed $\gamma>0$; for if not, we must have $h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)\right)=+\infty$ almost surely, which implies that
$\frac{1}{\eta_{*}}=\gamma_{*}^{-1}\mathbb{E}\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}\right)=\mathbb{E}\frac{1}{\gamma_{*}+h^{\prime\prime}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}\right)\right)}=0,$
contradicting Assumption 5. By this and Assumption 4, we know that Assumption 9 holds almost surely for all sufficiently large $p$. By Proposition C.12, almost surely for sufficiently large $p$, $g_{p}$ is well-defined, strictly increasing and equation (85) admits a unique solution on $(0,+\infty)$.
Now, (113), along with Assumption 1 and Proposition A.4, implies that almost surely
@@
Almost sure convergence (86) follows from (114), Proposition A.5 and the fact from Lemma C.13 that $1+\left(\mathsf{D}^{2}-\gamma\right)\mathbb{E}\frac{1}{\gamma+\mathsf{U}}>0$ almost surely. ∎
C.7 Finite or single coordinate inference under exchangeability
@@
We only show that $\frac{{r}_{*,i}-{\beta}^{\star}_{i}}{\sqrt{\tau_{*}}}\Rightarrow N(0,1)$ for $\mathbf{r}_{*}$ defined in (61). (29) then follows from consistency of $\hat{\tau}_{*}$ and $\widehat{\mathsf{adj}}$ (cf. Theorem C.2) and the Slutsky’s theorem. Let $\mathbf{U}\in\mathbb{R}^{p\times p}$ denote a permutation operator drawn uniformly at random independent of $\bm{\beta}^{\star},\mathbf{X},\bm{\varepsilon}$. We have that
@@
where we use $\stackrel{{\scriptstyle L}}{{=}}$ to denote equality in law. Note that
@@
$\qquad\stackrel{{\scriptstyle L}}{{=}}\left(\underset{\beta}{\operatorname{argmin}}\frac{1}{2}\left\|\mathbf{X}\left(\bm{\beta}^{\star}-\bm{\beta}\right)+\bm{\varepsilon}\right\|^{2}+h(\bm{\beta}),\mathbf{X},\bm{\beta}^{\star},\bm{\varepsilon}\right)$
@@
Below we prove the Corollary for $\mathcal{L}=\{i,k\},i\neq k$. The general case is analogous. For standard basis $\mathbf{e}_{i},\mathbf{e}_{k}$, and any constant $c_{1},c_{2}\in\mathbb{R}$,
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{P}\left(\frac{\mathbf{e}_{i}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{i}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{1},\frac{\mathbf{e}_{k}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{k}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{2}\right)$
$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{E}\left(\mathbb{P}\left(\frac{\mathbf{e}_{i}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{i}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{1},\frac{\mathbf{e}_{k}^{\top}\mathbf{U}\mathbf{r}_{*}-\mathbf{e}_{k}^{\top}\mathbf{U}\bm{\beta}^{\star}}{\sqrt{\tau_{*}}}<c_{2}\mid\mathcal{F}\left(\bm{\beta}^{\star},\bm{\varepsilon},\mathbf{X}\right)\right)\right)$
$\stackrel{{\scriptstyle(c)}}{{=}}\mathbb{E}\frac{1}{p(p-1)}\sum_{j_{1}\neq j_{2}\in[p]}\mathbb{I}\left(\frac{1}{\sqrt{\tau_{*}}}\left(r_{*,j_{1}}-\beta_{j_{1}}^{\star}\right)<c_{1}\right)\mathbb{I}\left(\frac{1}{\sqrt{\tau_{*}}}\left(r_{*,j_{2}}-\beta_{j_{2}}^{\star}\right)<c_{2}\right)$
@@
where $\mathsf{Z}\sim N(0,1)$. Here, we used Theorem C.2 and Proposition A.5. Using dominated convergence theorem, we conclude that
@@
C.8 Hypothesis testing and confidence intervals